Himani rolls two standard cubes. Find the probability that the sum of the roll is 5, given that one cube came up even and the other came up odd.

a) 1/36
b)1/9
c)5/36
d)2/9

Grade:12

2 Answers

bharat bajaj IIT Delhi
10 years ago
As it is given that one is even & one is odd, the sample space is not 36, it is only 18 !

Sample spacewith favourable outcomes :
1 & 4
4 & 1
2 & 3
3 & 2


Probability = 2/9
Thanks & Regards
Bharat Bajaj
askIITians Faculty
IIT Delhi
SREE NIGAMADITYA NIMMAGADDA
10 years ago
for this the outcomes can be 4&1, 1&4, 2&3, 3&2, but here we might think that sample space is `36` but its "18" since it is given that one is even and one is odd. so the probability would be equal to "4/18"=2/9